OpenStudy (anonymous):

28= 4/5 times Z + 5 z= ????

OpenStudy (anonymous):

28 = (4/5)z +5 23= (4/5)z z=23(5/4) = 115/4

OpenStudy (anonymous):

\[4/5 \times z + 5 \] I mean for 4/5 as a fraction

OpenStudy (anonymous):

28= (4/5)( Z + 5 ) 28*(5/4)-5 =z z=30